# Caiqi Zhang1, Fangyu Liu1, Marco Basaldella2, Nigel Collier1

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capability in a variety of NLP tasks. Despite their effectiveness, these models are prone to generate nonfactual content. Uncertainty Quantification (UQ) is pivotal in enhancing our understanding of a model’s confidence in its generated content, thereby aiding in the mitigation of nonfactual outputs. Existing research on UQ predominantly targets short text generation, typically yielding brief, word-limited responses. However, real-world applications frequently necessitate much longer responses. Our study first highlights the limitations of current UQ methods in handling long text generation. We then introduce Luq, a novel sampling-based UQ approach specifically designed for long text. Our findings reveal that Luq outperforms existing baseline methods in correlating with the model’s factuality scores (negative coefficient of -0.85 observed for Gemini Pro). With Luq as the tool for UQ, we investigate behavior patterns of several popular LLMs’ response confidence spectrum and how that interplays with the response’ factuality. We identify that LLMs lack confidence in generating long text for rare facts and a factually strong model (i.e. GPT-4) tends to reject questions it is not sure about. To further improve the factual accuracy of LLM responses, we propose a method called Luq-Ensemble that ensembles responses from multiple models and selects the response with the least uncertainty. The ensembling method greatly improves the response factuality upon the best standalone LLM.
@@ 1 Introduction
Although Uncertainty Quantification (UQ) is a well-researched area in the broader field of machine learning (Gawlikowski et al., 2023), its application in the context of LLMs remains relatively under-explored. One primary limitation is that previous studies on UQ mostly require access to a model’s internal states (e.g. logits) (Murray and Chiang, 2018; Kuhn et al., 2022; Vazhentsev et al., 2023; Duan et al., 2023). However, many of the most advanced LLMs, such as GPT-4 (OpenAI, 2023), Gemini Pro (Gemini Team, 2023), and Claude 2.1 (Anthropic, 2023), are closed-source and only accessible via API calls. This limits the ability to directly analyze their internal processes. Another challenge is that existing research on modeling uncertainty predominantly focuses on short responses, typically less than 10 words in length (Kuhn et al., 2022; Duan et al., 2023; Lin et al., 2023). This is in stark contrast to the more common use cases of LLMs, where responses to queries often far exceed this length, sometimes reaching hundreds of words. Such disparity points to a need for new UQ methods tailored for long-form text generated by LLMs. Therefore, in this study we aim to answer the following research questions: RQ1: Are existing UQ methods still effective in the context of long-text generation? RQ2: If not, how can we effectively quantify LLMs’ uncertainty for long-form answers? RQ3: In what ways can uncertainty scores be utilized to enhance the factuality of model outputs?
Our experiments primarily focus on black-box LLMs, with an emphasis on using factuality as the key metric to evaluate the models’ performance. The main contributions of this paper are:
  • We first highlight the limitations of existing UQ methods for long text generation and subsequently propose Luq (Long-text Uncertainty Quantification; pronounced as “luck”), a novel UQ method that computes sentence-level consistency in long text scenarios.
  • Through extensive experiments on the FActScore dataset, we demonstrate that our proposed Luq consistently shows strong negative correlations with the responses’ factuality over 6 popular LLMs, outperforming all the baseline methods.
@@ 2.1 Uncertainty and Confidence
Similarly, the concept of confidence within NLG frequently adopts methodologies from classification tasks. In these tasks, confidence for a specific prediction $y$ is quantified using the predicted probability, represented as $\hat{p}(Y=y\mid x)$ (Geifman and El-Yaniv, 2017; Hendrycks and Gimpel, 2016). In the context of NLG, the confidence score for a given response $\mathbf{r}$ is represented by the joint probability of the tokens in the response:
@@ 2.2 Uncertainty for Long Text Generation
There are two main limitations in applying traditional uncertainty and confidence definitions to long text generation. Firstly, while these definitions mostly rely on calculating predictive entropy to estimate the model’s uncertainty, this approach is not the only option and may not always be appropriate. Secondly, they depend heavily on accessing the model’s internal states, which can be difficult to obtain.
In our study, we adopt a more flexible definition of uncertainty and confidence. We move away from the assumption that confidence must directly reflect the model’s true accuracy in real-world scenarios—a benchmark that, particularly for long text generation, may often be unattainable. Instead, our approach places a premium on the model’s ability to rank samples effectively, distinguishing between correct and incorrect predictions.
From a practical perspective, we aim for the uncertainty score to serve as a reliable indicator of the model’s performance. This performance encompasses several dimensions of generation quality, including factuality, coherence, and creativity. Notably, our study prioritizes factuality and the truthfulness of responses, adopting these as our primary metrics. The factuality of the responses $\mathbf{R}$ given a specific query $x$ is denoted as $F\left(\mathbf{R}\mid\mathbf{x}\right)$. Considering two inputs $x_{i}$ and $x_{j}$, we explore the relationship between the model’s uncertainty, denoted as $\operatorname{U}\left(\mathbf{x}\right)$, and the factuality. Our goal is to have:
@@ 3 Luq
In this section, we introduce our Luq method for estimating uncertainty in long text generation. The framework is illustrated in Figure 1. Our underlying assumption posits that the greater the model’s uncertainty regarding a given question $x$, the more diverse its responses to question $x$ will be. For instance, as shown in Figure 1, the term “third pharaoh of the 20th Dynasty of Egypt” is frequently supported by other sample responses, indicating the model’s high confidence in this information. However, the samples suggest different reign periods for Ramesses IV; the inconsistency shows the model is uncertain about this information. We begin by highlighting the limitations of previous UQ methods on dealing with long text, then formally define our Luq method that overcomes these issues.
Motivation
Following the generation of $n$ responses, traditional UQ methods for short text commonly calculate the pairwise similarity among the responses, employing either Jaccard Similarity or calculating Natural Language Inference (NLI) scores (Kuhn et al., 2022; Lin et al., 2023). These pairwise similarity scores indicate the consistency between a pair of responses and play a vital role in subsequent uncertainty estimation.
However, responses to certain queries are brief, making it relatively straightforward to distinguish between different answers. For example, for the question “What is the capital of France?”, the answers often consists of just a few words and can be easily examined. In contrast, answers to other questions such as “Give me an introduction of …” and “Tell me something about …” may extend to hundreds of words. Longer text leads to an unexpected high similarity across all response pairs when applying previous methods. To address this issue and achieve a more nuanced similarity assessment, we propose the Luq uncertainty measurement with sentence-level similarity computation. Inspired by the hallucination detection method in Manakul et al. (2023), we split each response to sentences, and check whether each sentence can be supported by other samples.
Notation
Let $r_{a}$ represent the response generated by a LLM to a user query $x$. We generate an additional $n$ stochastic LLM sample responses $R=\{r_{1},r_{2},\ldots,r_{n}\}$ using the same query. The set $R^{\prime}=\{r_{a},r_{1},r_{2},\ldots,r_{n}\}$ encompasses all outputs from the model.
For any given response $r_{i}\in R^{\prime}$, the first objective is to determine how often it is supported (or entailed) by other samples. To this end, we employ a NLI classifier to assess the similarity between $r_{i}$ and each $r^{\prime}\in R^{\prime}\setminus\{r_{i}\}$. The output from an NLI classifier normally includes classifications of entailment, neutral, and contradiction, along with their respective logit values. It is important to note that we focus exclusively on the “entailment” and “contradiction” classes, as sentences labeled as “neutral” generally do not impact the overall factuality of a response. We calculate the NLI score for each sentence $s_{j}$ within a response $r$, and then average these scores. Formally, the similarity score $\mathcal{S}(r_{i},r^{\prime})$ between $r_{i}$ and $r^{\prime}$ is defined as:
@@
Unlike Kuhn et al. (2022)’s method of applying an off-the-shelf DeBERTa model, we apply the DeBERTa-v3-large model (He et al., 2022), fine-tuned on the MultiNLI (Williams et al., 2018) dataset. This choice is due to our input being a concatenation of short hypothesis (sentence $s$) and a comparatively longer premises (reference response $r^{\prime}$). The format of the input aligns with the task in MultiNLI dataset, ensuring an effective assessment of consistency among the responses.
@@ 4 Experiments
Methods GPT-4 GPT-3.5 Yi-34B-Chat Tulu-2-70B Gemini Pro Vicuna-33B
PCC SCC PCC SCC PCC SCC PCC SCC PCC SCC PCC SCC
LexSimilarity -45.25 -35.97 -67.75 -52.45 -70.09 -68.23 -55.69 -61.84 -67.24 -63.73 -38.28 -50.62
Eccentricity -24.83 -12.74 -10.61 -26.46 -27.63 -44.97 -2.13 10.12 -50.27 -57.79 -18.72 -13.98
NumSemSets -8.24 4.18 -11.90 -17.02 -25.69 -31.34 -20.68 -18.07 -52.98 -57.02 -20.04 -16.65
EigValLaplacian -36.94 -18.72 -30.30 -34.60 -48.97 -51.12 -50.08 -50.29 -72.73 -69.68 -60.49 -61.74
DegMat -3.78 6.73 -22.44 -22.85 -39.82 -38.94 -53.39 -53.97 -64.37 -67.69 -58.25 -62.38
SelfCheckNLI -53.12 -41.79 -65.14 -61.07 -70.34 -72.69 -75.63 -76.89 -84.49 -82.37 -66.82 -66.48
Luq -60.37 -45.28 -71.30 -66.62 -73.78 -74.59 -77.58 -75.39 -85.09 -81.29 -71.79 -70.78
Table 1: Pearson and Spearman correlation coefficients (expressed as percentages) between different LLMs and various UQ methods on the FactScore dataset. We use the original factuality scores instead of the penalized ones. All numbers are in percentages.
@@ 4.1 Dataset, Metric, and LLM Selection
Dataset
We use FActScore (Min et al., 2023) to assess the generated text’s factuality. We chose it as the dataset for two main reasons: firstly, we emphasize the critical importance of factuality in text generation, arguing that LLMs generated text should prioritize truthfulness. We hypothesize that increased uncertainty in a model correlates with the production of factually inaccurate information. Secondly, FActScore allows for automated evaluation with a low error rate (below 2%), making it scalable for applying to a wide array of new LLMs without human annotation. The dataset includes 500 individuals from Wikidata with corresponding Wikipedia entries.
Metrics
For each generated response, FActScore calculates a factuality score (FS). It first breaks down the response into a series of atomic facts, which are concise statements that capture a single piece of information. It then assigns a binary label to each atomic fact and calculate the precision as the response’s factuality score. We apply FActScore for the first generated response ($r_{a}$). As the LLMs may sometime refuse to answer certain questions, to have a fair comparison, we introduce a penalized factuality score (PFS) and penalized uncertainty score (PUS). To calculate PFS and PUS, we assign a factuality score of zero and uncertainty score of one to questions that models opt not to answer.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capability in a variety of NLP tasks. However, LLMs are also prone to generate nonfactual content. Uncertainty Quantification (UQ) is pivotal in enhancing our understanding of a model’s confidence on its generation, thereby aiding in the mitigation of nonfactual outputs. Existing research on UQ predominantly targets short text generation, typically yielding brief, word-limited responses. However, real-world applications frequently necessitate much longer responses. Our study first highlights the limitations of current UQ methods in handling long text generation. We then introduce Luq and its two variations, a series of novel sampling-based UQ approaches specifically designed for long text. Our findings reveal that Luq outperforms existing baseline methods in correlating with the model’s factuality scores (negative coefficient of -0.85 observed for Gemini Pro). To further improve the factuality of LLM responses, we propose Luq-Ensemble, a method that ensembles responses from multiple models and selects the response with the lowest uncertainty. The ensembling method greatly improves the response factuality upon the best standalone LLM.
@@ 1 Introduction
Although Uncertainty Quantification (UQ) is a well-researched area in machine learning (Gawlikowski et al., 2023), its application in the context of LLMs remains under-explored. One primary limitation is that previous studies on UQ mostly require access to a model’s internal states (e.g. logits) (Murray and Chiang, 2018; Kuhn et al., 2022; Vazhentsev et al., 2023; Duan et al., 2023). However, many best-performing LLMs, such as GPT-4 (OpenAI, 2023), Gemini 1.0 Pro (Gemini Team, 2023), and Claude 2.1 (Anthropic, 2023), are closed-source and only accessible via API calls. This limits the ability to directly analyze their internal processes. Another challenge is that existing research on modeling uncertainty predominantly focuses on short responses, typically less than 10 words in length (Kuhn et al., 2022; Duan et al., 2023; Lin et al., 2023). This is in stark contrast to the more common use cases of LLMs, where responses to queries often far exceed this length, sometimes reaching hundreds of words. Such disparity points to a need for new UQ methods tailored for long-form text generated by LLMs. Therefore, in this study we aim to answer the following research questions: RQ1: Are existing UQ methods still effective in the context of long-text generation? RQ2: If not, how can we effectively quantify LLMs’ uncertainty for long-form answers? RQ3: In what ways can uncertainty scores be utilized to enhance the factuality of model outputs?
We explore UQ for long-text generation (at least 100 words), with an emphasis on using factuality as the key metric of the models’ performance. The main contributions of this paper are:
  • We first highlight the limitations of existing UQ methods for long text generation and then propose Luq (Long-text Uncertainty Quantification; pronounced as “luck”), a novel UQ method that computes sentence-level consistency in long text scenarios.
  • Through extensive experiments on the original FActScore dataset and our newly proposed FActScore-Dis dataset in medical domain, we demonstrate that Luq consistently shows strong negative correlations with the responses’ factuality over 6 popular LLMs, outperforming all the baseline methods.
@@ 2.1 Uncertainty and Confidence
Currently, a formal and universally accepted definition of uncertainty levels in language generation tasks remains elusive. Common practice in existing literature measures uncertainty through the entropy of predictions, akin to approaches in classification tasks (Kuhn et al., 2022; Lin et al., 2023). Predictive entropy is formally expressed as:
@@
It captures the uncertainty associated with a prediction for a given input $x$. In the context of NLG, where $\mathbf{R}$ denotes all possible generations and $\mathbf{r}$ is a specific response, the uncertainty score can thus be conceptualized as:
@@
In classification tasks, confidence for a specific prediction $y$ is quantified using the predicted probability, represented as $\hat{p}(Y=y\mid x)$ (Geifman and El-Yaniv, 2017; Hendrycks and Gimpel, 2016). Similarly, in the context of NLG, the confidence score for a given response $\mathbf{r}$ is represented by the joint probability of the tokens in the response:
@@ 2.2 Uncertainty for Long Text Generation
In our study, we adopt a more flexible approach to defining uncertainty and confidence in long text generation. Similar to Huang et al. (2024), we focus on the ability of UQ methods to effectively rank responses, differentiating between correct and incorrect predictions. This approach also aligns with the concept of relative confidence as discussed by Geng et al. (2023). Our objective diverges from the orthogonal research direction about models’ calibration, which requires models to precisely reflect their true accuracy in practical scenarios (Lin et al., 2023). We argue that while short-answer questions may be straightforwardly assessed using metrics such as accuracy or exact match, these standards are often unrealistic for long text generation, given the complexities of real-life probabilities.
From a practical perspective, we aim for the uncertainty score to serve as a reliable indicator of the model’s performance. This performance encompasses several dimensions of generation quality, including factuality, coherence, and creativity. Our study prioritizes factuality and the truthfulness of responses, adopting these as our primary metrics. The factuality of the responses $\mathbf{R}$ given a specific query $x$ is denoted as $F\left(\mathbf{R}\mid\mathbf{x}\right)$. Considering two inputs $x_{i}$ and $x_{j}$, we explore the relationship between the model’s uncertainty, denoted as $\operatorname{U}\left(\mathbf{x}\right)$, and the factuality. Our goal is to have:
@@ 3 Luq
In this section, we introduce our Luq method and its two variations (Luq-Atomic and Luq-Pair) to estimate uncertainty in long text generation. The overall framework is illustrated in Figure 1.
Motivation. Our underlying assumption posits that the greater the model’s uncertainty regarding a given question $x$, the more diverse its responses to question $x$ will be. For instance, as shown in Figure 1, the term “third pharaoh of the 20th Dynasty of Egypt” is frequently supported by other sample responses, indicating the model’s high confidence in this information. However, the samples suggest different reign periods for Ramesses IV; the inconsistency shows the model’s higher uncertainty.
Following the generation of $n$ responses, traditional UQ methods for short text commonly calculate the pairwise similarity among the responses (Kuhn et al., 2022; Lin et al., 2023). These pairwise similarity scores indicate the consistency between a pair of responses and play a vital role in subsequent uncertainty estimation. However, answers to certain questions such as “Give me an introduction of …” and “Tell me something about …” may extend to hundreds of words. Longer text leads to an unexpected high similarity across all response pairs when applying previous methods. To address this issue and achieve a more nuanced similarity assessment, we propose the Luq uncertainty measurement with sentence-level similarity computation. Inspired by the hallucination detection method in Manakul et al. (2023), we split each response to sentences, and check whether each sentence can be supported by other samples.
Notation. Let $r_{a}$ represent the response generated by a LLM to a user query $x$. We generate an additional $n$ stochastic LLM sample responses $R=\{r_{1},r_{2},\ldots,r_{n}\}$ using the same query. The set $R^{\prime}=\{r_{a},r_{1},r_{2},\ldots,r_{n}\}$ encompasses all outputs from the model.
For any given response $r_{i}\in R^{\prime}$, the first objective is to determine how often it is supported (or entailed) by other samples. To this end, we employ an NLI classifier to assess the similarity between $r_{i}$ and each $r^{\prime}\in R^{\prime}\setminus\{r_{i}\}$. The output from an NLI classifier normally includes classifications of entailment, neutral, and contradiction, along with their respective logit values. It is important to note that we focus exclusively on the “entailment” and “contradiction” classes, as sentences labeled as “neutral” generally do not impact the overall factuality of a response. We calculate the NLI score for each sentence $s_{j}$ within a response $r$, and then average these scores. Formally, the similarity score $\mathcal{S}(r_{i},r^{\prime})$ between $r_{i}$ and $r^{\prime}$ is defined as:
@@
Unlike Kuhn et al. (2022)’s method of applying an off-the-shelf DeBERTa model, we apply the DeBERTa-v3-large model (He et al., 2022), fine-tuned on the MultiNLI (Williams et al., 2018) dataset. This choice is due to our input being a concatenation of short hypothesis (sentence $s$) and a comparatively longer premises (reference response $r^{\prime}$). The format of our input aligns with the task in MultiNLI dataset, ensuring an effective assessment of consistency among the responses.
Luq-Atomic. To check the consistency of the generated responses in a more fine-grained manner, we implement Luq-Atomic, a variation of the original Luq. The key difference is that it first uses ChatGPT to break a response $r$ into atomic fact pieces $\{a_{1},a_{2},...,a_{j}\}$. Luq-Atomic then calculates the uncertainty scores bases on atomic fact piece level ($a_{j}$) instead of sentence level ($s_{j}$).
Luq-Pair. The performance of our NLI classifier may be constrained by the length of the premises and hypotheses. To address this, we propose Luq-Pair to calculate the entailment score $s_{j}$ for each sentence $s^{\prime}_{j}$ in $r^{\prime}$ and select the maximum value. Formally, we define this as:
$\mathcal{P}\left(\text{entail}\mid s_{j},r^{\prime}\right)=\max_{s^{\prime}_{j}\in r^{\prime}}\left|\mathcal{P}\left(\text{entail}\mid s_{j},s^{\prime}_{j}\right)\right|$
We discuss more about Luq-Atomic and Luq-Pair in Appendix A.
@@ 4 Experiments
White-Box Methods Black-Box Methods
MSP MCSE SE LexSim Ecc NumSets EigV Deg SCN Luq
FActScore-Bio
GPT-4 PCC - - - -45.2 -24.8 -8.24 -36.9 -3.78 -53.1 -60.4
SCC - - - -36.0 -12.7 4.18 -18.7 6.73 -41.8 -45.3
GPT-3.5 PCC - - - -67.8 -10.6 -11.9 -30.3 -22.4 -65.1 -71.3
SCC - - - -52.4 -26.5 -17.0 -34.6 -22.9 -61.1 -66.6
Gemini 1.0 Pro PCC - - - -67.2 -50.3 -53.0 -72.7 -64.4 -84.5 -85.1
SCC - - - -63.7 -57.8 -57.0 -69.7 -67.7 -82.4 -81.3
Yi-34B-Chat PCC -20.7 -43.9 -55.8 -70.1 -27.6 -25.7 -49.0 -39.8 -70.3 -73.8
SCC -22.1 -44.3 -53.6 -68.2 -45.0 -31.3 -51.1 -38.9 -72.7 -74.6
Tulu-2-70B PCC -16.8 -32.4 -50.5 -55.7 -2.13 -20.7 -50.1 -53.4 -75.6 -77.6
SCC -15.4 -34.8 -52.7 -61.8 10.1 -18.1 -50.3 -54.0 -76.9 -75.4
Vicuna-33B PCC -28.5 -36.8 -58.6 -38.3 -18.7 -20.0 -60.5 -58.3 -66.8 -71.8
SCC -27.9 -37.4 -57.2 -50.6 -14.0 -16.6 -61.7 -62.4 -66.5 -70.8
FActScore-Dis
GPT-3.5 PCC - - - -41.8 -27.9 -7.81 -38.8 -13.5 -59.0 -67.3
SCC - - - -39.4 -26.0 -6.94 -36.9 -16.3 -59.1 -65.3
Yi-34B-Chat PCC -20.3 -35.4 -52.6 -63.6 -19.3 -11.2 -40.6 -26.5 -65.1 -70.5
SCC -21.7 -33.8 -54.9 -58.7 -21.5 -16.3 -38.4 -22.1 -67.8 -72.4
Table 1: Pearson and Spearman correlation coefficients (expressed as percentages) between different LLMs and various UQ methods on the FactScore dataset. We use the original factuality scores instead of the penalized ones.
@@ 4.1 Dataset, Metric, and LLM Selection
Dataset. When selecting the dataset, we considered three main criteria: (1) The dataset should be a long-form QA dataset. (2) There should be a well-designed and widely-accepted automatic evaluation tool. (3) The questions should be clear, specific, and have definite answers for objective evaluation. We therefore employ FActScore (Min et al., 2023) to evaluate the factuality of our generated text. It offers automated assessment with a low error rate (below 2%), enabling scalable application to diverse LLMs without requiring manual annotation. To supplement the extensive reliability testing of FActScore conducted by its creators, we performed a smaller-scale human annotation study. Our findings demonstrate a strong Pearson correlation of 0.88 between FActScore ratings and human factuality judgments, confirming it being a reliable reference for factuality. Please refer to Appendix B for more information about the discussion of this dataset and our validation process.
The original FActScore dataset (denoted as FActScore-Bio) includes 500 individuals’ biographies from Wikidata with corresponding Wikipedia entries. To evaluate the applicability of UQ methods across different domains, we additionally developed a dataset, FActScore-Dis, focusing on disease entities. Details of this dataset can be found in Appendix C.
Metrics. For each generated response, FActScore calculates a factuality score (FS). We apply FActScore for the first generated response ($r_{a}$). As the LLMs may sometime refuse to answer certain questions, to have a fair comparison, we introduce a penalized factuality score (PFS) and penalized uncertainty score (PUS). To calculate PFS and PUS, we assign a factuality score of zero and uncertainty score of one to questions that models opt not to answer.